OpenStudy (anonymous):

i need to express x/1-x as a power series

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry bout that mate

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so the series is \[\sum_{0}^{\infty} x / (1-x)\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

im sorry that is completely wrong. 1/(1-x) = \[\sum_{0}^{\infty} x^n\] im sorry that is completely wrong

OpenStudy (anonymous):

where you are sending 0 --> infinity

OpenStudy (anonymous):

what a series does is it models something else by using and infinite amount of terms. that is why you can equate what i stated above

OpenStudy (anonymous):

and you sum allll of those terms

OpenStudy (anonymous):

1/(1-x) is defined by \[\sum_{n \rightarrow0}^{\infty}x^n\] = \[1 + x^2 + x^3 + x^4 + x^5 + ... + x^n\] but this is not your equation.

OpenStudy (anonymous):

right? you can find that series in any caluculus text book

OpenStudy (anonymous):

that series has to be memorized

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so to get to your equation you must multiply everyyy term by x because x * 1/(1-x) = \[x *\sum_{n=0}^{\infty} x^n\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

multiplying by x your exponents are n and 1. when multiplying like bases you add exponents you get n + 1 so. \[x/(1-x) = \sum_{n=0}^{\infty} x^{n+1}\] = \[x + x^2 + x^3 + x^4 + x^5 + .... + x^{n+1}\]

OpenStudy (anonymous):

cool right?

OpenStudy (anonymous):

ok can i ask you one more question

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so to find the interval of convergence how would i go about that

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

have you learned the nth root test? or ratio test?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so you can take the ratio test. of the series. which states you must take the absolute value of the An+1 and divide by An you should get | x^[ (n+1) +1 ] / x^(n+1) | = | x^(n+2) / x^(n+1) |

OpenStudy (anonymous):

can you simplify that? dont froget the ABS and show me what you get

OpenStudy (anonymous):

Is that just X^N?

OpenStudy (anonymous):

|x^N|

OpenStudy (anonymous):

you can look at it as multiples in the numerator and denominator

OpenStudy (anonymous):

you have x^(n+2) = x^n * x * x right? its all just boring exponent work

OpenStudy (anonymous):

you get that?

OpenStudy (anonymous):

in the numerator

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you get that?

OpenStudy (anonymous):

So is it just X^N+1?

OpenStudy (anonymous):

dont forget you have a multiple of x^n in the denominator so they cancel each other x^(n+2) / x^(n+1) = { x^n * x * x } / { x^n * x }

OpenStudy (anonymous):

So just X then?

OpenStudy (anonymous):

yeah. real quick work it out on a piece of paper while i type up the next part

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the ratio test states Xn+1 / Xn < \[\left| x_{n+1}/x _{n} \right|< \rho =1\] Xn+1 / Xn = x

OpenStudy (anonymous):

you can plug solve for x in the inequality | x | < p = 1 | x | < 1 -1 > x > 1

OpenStudy (anonymous):

so that is the IOC

OpenStudy (anonymous):

ah ok got it

OpenStudy (anonymous):

you can test the end points now An = X^(n+1)

OpenStudy (anonymous):

so to check for the end points you just plug into x

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

and for 1 it diverges because it doesnt converge and -1 it diverges also right

OpenStudy (anonymous):

but into the series. its kind of something a little technical x^n doesnt equal series x^n

OpenStudy (anonymous):

yeah because it is a divergent geometric series where r must be <1 r^n

OpenStudy (anonymous):

or acutally its |r| < 1 to converge and your r's are -1 and 1. any x that is a fraction, for this series, would converge.

OpenStudy (anonymous):

alright got it

OpenStudy (anonymous):

the ratio test is just difficult with the exponents

OpenStudy (anonymous):

ok thanks